# When Respondents Don’t Care Anymore: Identifying the Onset of Careless Responding

### v3

@@ 7 Initialize $\{\textsf{LSP}_{j}\}_{j=1}^{p}$;
B.4 Details on Changepoint Detection
Let $\{\bm{Y}_{j}\}_{j=1}^{p}$ be a series of length $p$ consisting of $d$-dimensional random variables. The goal of the cumulative sum self-normalization test of Shao & Zhang, (2010) is to estimate the location of a possible changepoint in the (multivariate) series, i.e., to divide the series into two subsets $\{\bm{Y}_{j}\}_{j=1}^{k-1}$ and $\{\bm{Y}_{j}\}_{j=k}^{p}$ in which a $d$-dimensional parameter of interest $\bm{\theta}$ differs. Here, we focus on the mean as the parameter $\bm{\theta}$. Define by $\widehat{\bm{\theta}}_{a,b}=(b-a+1)^{-1}\sum_{j=a}^{b}\bm{Y}_{j}$ the mean of the series calculated on the subset implied by periods $a\leq b$. For location $k\in\{2,\dots,p\}$, define the statistic
@@ B.4 Details on Changepoint Detection
$\bm{L}_{p}(k)=\left\{\begin{array}[]{ll}\displaystyle\sum_{i=1}^{k-2}\frac{i^{2}(k-1-i)^{2}}{p^{2}(k-1)^{2}}\left(\widehat{\bm{\theta}}_{1,i}-\widehat{\bm{\theta}}_{i+1,k-1}\right)\left(\widehat{\bm{\theta}}_{1,i}-\widehat{\bm{\theta}}_{i+1,k-1}\right)^{\top},&\qquad\text{if }k>2,\\ 0,&\qquad\text{if }k=2,\\ \end{array}\right.$
@@
$\bm{R}_{p}(k)=\left\{\begin{array}[]{ll}\displaystyle\sum_{i=k+1}^{p}\frac{(p-i+1)^{2}(i-k)^{2}}{p^{2}(k-1)^{2}}\left(\widehat{\bm{\theta}}_{i,p}-\widehat{\bm{\theta}}_{k,i-1}\right)\left(\widehat{\bm{\theta}}_{i,p}-\widehat{\bm{\theta}}_{k,i-1}\right)^{\top},&\qquad\text{if }k<p,\\ 0,&\qquad\text{if }k=p,\\ \end{array}\right.$
@@
Otherwise, no changepoint is detected. Shao & Zhang, (2010) and Zhao et al., (2022) derive theoretical guarantees for this procedure. Specifically, Shao & Zhang, (2010) obtain the asymptotic distribution of the test statistic $SN_{p}$ for $p\rightarrow\infty$ and provide the corresponding critical values $K_{d}$ for a given significance level $\alpha$ and dimension $d$. We summarize relevant critical values in Table B.2.
Table B.2: Critical values $K_{d}$ for relevant significance levels $\alpha$ and dimensions $d$.
@@
Note. The critical values are taken from Table 1 in Shao & Zhang, (2010).
In the context of this paper, we use the test statistic from (B.3) to test for changepoints in the two-dimensional series
@@
where $\textsf{RE}_{j}$ is the autoencoder reconstruction error from (1) of a response to item $j$, and $\textsf{LSP}_{j}$ is the value of the LSP sequence from (2) assigned to this response. If a changepoint $\widehat{k}$ is detected, under Assumptions 4 and 5, the subsets $\{\bm{Y}_{j}\}_{j=1}^{\widehat{k}-1}$ and $\{\bm{Y}_{j}\}_{j=\widehat{k}}^{p}$ may be interpreted as attentive and careless responses, respectively. That is, the detected changepoint $\widehat{k}$ may be viewed as the onset of careless responding. We obtain such a series for each of the $n$ survey participants and subsequently test each of the $n$ series for changepoints.
@@ Appendix C Additional Simulations
C.1 Additional Results for the Main Design
In addition to CODERS, we apply three traditional methods for the detection of careless respondents in the simulation design from the main text, which is inspired by the data collected by Johnson, (2005): personal reliability, psychometric antonym, and the longstring index. We choose these three methods since they were also applied in the study of Johnson, (2005). Following Johnson,, we reverse psychometric antonym scores, and we flag respondents as careless whose personal reliability and psychometric antonym scores are lower than $0.3$ and $-0.03$, respectively. For the longstring index, we flag a respondent as careless if its value exceeds 6 (i.e., $>6$), which is a cutoff value suggested by Niessen et al., (2016).
Figure C.1 visualizes the false positive rate (FPR) and the false negative rate (FNR) of the three methods for the different carelessness onset regimes (in separate columns). Figure C.2 further shows the false negative rates for the four types of partially careless respondents (in separate rows). The longstring index performs perfectly at its intended task of detecting straightlining, whereas it hardly detects any random responding or pattern responding (which is expected). Furthermore, it performs better than expected at identifying extreme responding. Since the survey is lengthy and those respondents choose randomly from only the two extreme answer categories, there is a relatively high probability (depending on the onset) that they select a long string of the same answer category somewhere by chance. It turns out that personal reliability flags hardly any respondents, resulting in false positive rates close to 0 and false negative rates close to 1. Psychometric antonym yields inflated false positive rates, while the false negative rates are still relatively high. Interestingly, its performance is fairly constant across the different types of carelessness. As expected, the false negative rate of the traditional methods in general worsens the later the onset of careless responding (while that of CODERS deteriorates the earlier the onset, see Figures 7 and 8 in the main text).
Figure C.1: Performance measures for the traditional methods used by Johnson, (2005).
@@ C.1 Additional Results for the Main Design
Figure C.2: False negative rate of the traditional methods used by Johnson, (2005) for different types of partially careless respondents.
@@
C.2 Decreasing Survey Length
Recall that the changepoint detection method of Shao & Zhang, (2010) used in CODERS is based on asymptotic arguments for the number of items being large (see Appendix B.4). It therefore seems plausible that the ability of CODERS to accurately identify partial carelessness may decline as survey length decreases.
C.2.1 Data Generation
@@
C.2.2 Results
Figure C.3 plots the false positive rates for the different survey lengths (in separate columns). As before, CODERS rarely incorrectly flags attentive respondents as careless at any of the considered significance levels, although there is some overrejection for the shortest survey with $p=60$. Figure C.4 visualizes the false negative rate (top row) and the mean absolute error of onset estimation (bottom row) for careless respondents across the different surveys lengths (in separate columns). At a given significance level, the false negative rate slightly increases with decreasing survey length compared to the results from the main text. For $p=60$, the false negative rate reaches about 30%–40%, depending on the significance level. In addition, the location of the onset of carelessness is still estimated highly accurately throughout the investigated survey lengths, on average being off by only about one or two items.
Figure C.3: False positive rate of CODERS for attentive respondents across decreasing survey length.
@@ C.2.2 Results
Figure C.4: Performance measures of CODERS for partially careless respondents across decreasing survey length.
@@
Figure C.5: False negative rate of CODERS for different types of partially careless respondents across decreasing survey length.
@@
Figure C.6: Mean absolute error of carelessness onset estimation with CODERS for different types of partially careless respondents across decreasing survey length.
@@
Figures C.5 and C.6 show the false negative rate and the mean absolute error of onset estimation, respectively, grouped by carelessness type (rows) for each considered survey length (columns). In essence, the slight increase in false negative rates seems to be primarily caused by fewer inconsistent careless respondents being detected, whereas the performance for invariable careless responding stays near-perfect. For $p=60$, in particular very few random respondents are correctly identified, but the performance is still acceptable for extreme respondents.
@@
C.3 Additional Sample Sizes
@@
C.3.1 Data Generation
[…]
C.3.2 Results
Figures C.7C.10 visualize the performance measures for $n=248$ and $n=1,000$ simulated participants (in separate columns). The results are similar to those in the main text, with the overall FNR in partially careless respondents increasing only slightly for the smaller sample size and decreasing only slightly for the larger sample size. For random respondents, however, the increase in false negatives in the smaller sample size is notable, but so is the decrease in FNR in the larger sample size (see Figure C.9, top row). This can likely be explained by the amount of information the autoencoder has available in order to learn to distinguish between attentive responding and random responding. It is therefore conceivable that the power for detecting random respondents may improve further in even larger sample sizes.
Figure C.7: False positive rate of CODERS for attentive respondents across additional sample sizes.
@@ C.3.2 Results
Figure C.8: Performance measures of CODERS for partially careless respondents across additional sample sizes.
@@
Figure C.9: False negative rate of CODERS for different types of partially careless respondents across additional sample sizes.
@@
Figure C.10: Mean absolute error of carelessness onset estimation with CODERS for different types of partially careless respondents across additional sample sizes.
@@
C.4 Randomly Choosing from the Middle Answer Categories
@@ C.4.1 Data Generation
C.4.2 Results
Figures C.11 and C.12 visualizes the performance of CODERS when partially careless respondents are of the middling type. One the one hand, CODERS still almost never incorrectly flags an attentive respondent as partially careless. On the other hand, it identifies very few middling respondents with a false negative rate between about 85% and 95%, depending on the significance level. Hence, these simulations further demonstrate that CODERS requires inconsistent responding to be sufficiently different from attentive responding in order to distinguish between the two.
Figure C.11: False positive rate of CODERS for attentive respondents in the presence of partially careless respondents that randomly choose from the middle answer categories.
@@ C.4.2 Results
Figure C.12: Performance measures of CODERS for partially careless respondents that randomly choose from the middle answer categories.
@@
In order to further highlight the difficulty of detecting middling carelessness, we again apply the traditional methods personal reliability, psychometric antonym, and the longstring index. The false positive rates and false negative rates are shown in Figure C.13. As in the simulation design from the main text, personal reliability flags hardly any respondents, resulting in false positive rates close to 0 and false negative rates close to 1. The performance of the longstring index is similar, which is not surprising because it was designed for detecting straightlining (invariable carelessness) rather than inconsistent responding like middling. Psychometric antonym identifies more middling respondents than CODERS with a false negative rate of about 65%–75%, but it exhibits an inflated false positive rate of up to about 15%.
Figure C.13: Performance measures of the traditional methods for partially careless respondents that randomly choose from the middle answer categories.
@@ C.5.1 Data Generation
C.5.2 Results
Figures C.14 and C.15 visualize the performance of CODERS and Figures C.16 and C.17 that of the traditional methods. Clearly, CODERS rarely flags any participants, with a false positive rate close to 0 and a false negative rate close to 1 (we therefore omit separate results for the four types of partially careless responding). The same behavior is observed for personal reliability. Psychometric antonym achieves a better false negative rate (about 70%–85%) but struggles with many false positives (up to about 15%). The longstring index yields somewhat better performance, with a relatively low false positive rate and a false negative rate around 65%. As in the simulation design from the main text, it successfully captures straightliners as well as some extreme respondents (see Appendix C.1 for a more detailed discussion). In summary, CODERS is simply not designed for detecting more than one changepoint in respondent behavior, such that an extension to multiple changepoint detection may be a fruitful direction for further research.
Figure C.14: False positive rate of CODERS for attentive respondents in case of temporary carelessness.
@@ C.5.2 Results
Figure C.15: Performance measures of CODERS for partially careless respondents in case of temporary carelessness.
@@
Figure C.16: Performance measures for the traditional methods in case of temporary carelessness.
@@
Figure C.17: False negative rate of the traditional methods for different types of partially careless respondents in case of temporary carelessness.
@@
C.6 Organizational Survey
Until here, all simulations are based on a measurement of the Big Five personality traits. To study how CODERS fares in surveys of different nature, we further apply a simulation design based on a study from organizational research.
C.6.1 Data Generation
We simulate a hypothetical survey inspired by a study conducted among university students working in teams on a project for a Strategic Management course. The measurement includes various individual- and team-level constructs, among others the HEXACO-60 personality assessment of Ashton & Lee, (2009), the goal orientation scales of Vandewalle, (1997), the transactive memory systems scales of Lewis, (2003), the intra-group conflict scales of Jehn, (1995), the intra-unit justice climate scales of Li & Cropanzano, (2009), a team commitment scale based on Mowday et al., (1979), and the team performance scale of Hackman, (1986). In total, the hypothetical survey includes $p=198$ items measuring $q=34$ constructs (which we set as the number of nodes in the bottleneck layer of the autoencoder in CODERS).
Figure C.18: Correlation matrix and marginal response distributions used in the simulated organizational survey.
(a) Correlation matrix of the $p=198$ items. A more saturated blue (red) in a cell corresponds to a stronger positive (negative) correlation between the corresponding items.
(b) Marginal response distributions of the $p=198$ items ($x$-axis), each of which has five response options ($y$-axis). A darker shade of gray indicates a higher probability of the corresponding response option.
Note. The correlation matrix and the marginal response distributions are calculated from empirical data. Of the $p=198$ items, 152 are positively scored and 46 are negatively scored. The items in both plots are presented in an order such that items measuring the same construct are adjacent.
Although team-level constructs are measured, we simplify the data generating process and simulate individual-level data using the correlation matrix and marginal distributions displayed in Figure C.18, which are based on empirical data from 322 students (after removing 29 respondents that contain missing values). Unfortunately, we are not at liberty to share the original data. Of the $p=198$ items, 152 are positively scored and 46 are negatively scored. The different constructs are clearly visible along the diagonal of the correlation matrix, as are various relationships between many of the measured constructs in the off-diagonal parts (cf. also Alfons et al., 2022a, ; Alfons et al., 2022b, ). We randomize the order of the items in a way that ensures that items measuring the same construct remain adjacent, which is better suited for surveys in organizational contexts. That is, we first randomize the order of the items within each construct, and we randomize the order of (groups of related) constructs.
Otherwise, we keep the same design as in the simulations from the main text. In particular, we simulate responses for $n=500$ participants, using the same settings for generating partially careless respondents.
C.6.2 Results
Figure C.19 shows the false positive rates for attentive respondents. Compared to the simulations from the main text, false positive rates are inflated for high prevalence levels, but they remain reasonably close to the nominal level while carelessness prevalence does not exceed 40% (in the early onset regime) or 30% (for the other onset regimes). A possible explanation for this finding may be that many measured constructs do not use negatively-scored items while others do, such that response behavior may be naturally different in some parts of the survey than in others.
Figure C.19: False positive rate of CODERS for attentive respondents in the simulated organizational survey.
In Figures C.20C.22, the performance measures for (different types) of partially careless respondents are displayed. The overall false negative rate is similar to that of the simulations from the main text, despite the shorter survey length. But interestingly, there are slightly fewer false negatives for random responding and slightly more false negatives for extreme responding. This may be explained by differences in response distributions, with higher skewness in some of the measurements compared to the data by Johnson, (2005) (cf. Figure C.18 with Figure 5 from the main text). In addition, the onset of careless responding is estimated with similar accuracy, with an error of about two to three items averaged across the different types of partially careless responding. All in all, these results are an indication that the performance of CODERS may not depend too much on the institutional setting of the survey.
Figure C.20: Performance measures of CODERS for partially careless respondents in the simulated organizational survey.
Figure C.21: False negative rate of CODERS for different types of partially careless respondents in the simulated organizational survey.
Figure C.22: Mean absolute error of carelessness onset estimation with CODERS for different types of partially careless respondents in the simulated organizational survey.

### v4

@@ Appendix B Details on CODERS
Details on Changepoint Detection
Let $\{\bm{Y}_{j}\}_{j=1}^{p}$ be a series of length $p$ consisting of $d$-dimensional random variables. The goal of the cumulative sum self-normalization test of Shao and Zhang (2010) is to estimate the location of a possible changepoint $k$ in the (multivariate) series, i.e., to divide the series into two subsets $\{\bm{Y}_{j}\}_{j=1}^{k-1}$ and $\{\bm{Y}_{j}\}_{j=k}^{p}$ in which a $d$-dimensional parameter of interest $\bm{\theta}$ differs. Here, we focus on the mean as the parameter $\bm{\theta}$. Define by $\widehat{\bm{\theta}}_{a,b}=(b-a+1)^{-1}\sum_{j=a}^{b}\bm{Y}_{j}$ the mean of the series calculated on the subset implied by periods $a\leq b$. For location $k\in\{2,\dots,p\}$, define the statistic
@@
[…]
@@
[…]
@@
Otherwise, no changepoint is detected. Shao and Zhang (2010) and Zhao et al. (2022) derive theoretical guarantees for this procedure. Specifically, Shao and Zhang (2010) obtain the asymptotic distribution of the test statistic $SN_{p}$ for $p\rightarrow\infty$ and provide the corresponding critical values $K_{d}$ for a given significance level $\alpha$ and dimension $d$. We summarize relevant critical values in Table B2.
Table B2: Critical values $K_{d}$ for relevant significance levels $\alpha$ and dimensions $d$.
@@
Note. The critical values are taken from Table 1 in Shao and Zhang (2010).
In the context of this paper, we use the test statistic from (B3) to test for changepoints in the two-dimensional series
@@
where $\textsf{RE}_{j}$ is the autoencoder reconstruction error from (1) of a response to item $j$, and $\textsf{LSP}_{j}$ is the value of the LSP sequence from (2) assigned to this response. If a changepoint $\widehat{k}$ is detected, under Assumptions 4 and 5, the subsets $\{\bm{Y}_{j}\}_{j=1}^{\widehat{k}-1}$ and $\{\bm{Y}_{j}\}_{j=\widehat{k}}^{p}$ may be interpreted as attentive and careless responses, respectively. That is, the detected changepoint $\widehat{k}$ may be viewed as the onset of careless responding. We obtain a series $\{\bm{Y}_{j}\}_{j=1}^{p}$ for each of the $n$ survey participants and subsequently test each of the $n$ series for changepoints.
@@ Appendix C Additional Simulations
Additional Results for the Main Design
In addition to CODERS, we apply three traditional methods for the detection of careless respondents in the simulation design from the main text, which is inspired by the data collected by Johnson (2005): personal reliability, psychometric antonym, and the longstring index. We choose these three methods since they were also applied in the study of Johnson (2005). Following Johnson, we reverse psychometric antonym scores, and we flag respondents as careless whose personal reliability and psychometric antonym scores are lower than $0.3$ and 0.03-0.03, respectively. For the longstring index, we flag a respondent as careless if its value exceeds 6 (i.e., >6>6), which is a cutoff value suggested by Niessen et al. (2016).
Figure C1 visualizes the false positive rate (FPR) and the false negative rate (FNR) of the three methods for the different carelessness onset regimes (in separate columns). Figure C2 further shows the false negative rates for the four types of partially careless respondents (in separate rows). The longstring index performs perfectly at its intended task of detecting straightlining, whereas it hardly detects any random responding or pattern responding (which is expected). Furthermore, it performs better than expected at identifying extreme responding. Since the survey is lengthy and those respondents choose randomly from only the two extreme answer categories, there is a relatively high probability (depending on the onset) that they select a long string of the same answer category somewhere by chance. It turns out that personal reliability flags hardly any respondents, resulting in false positive rates close to 0 and false negative rates close to 1. Psychometric antonym yields inflated false positive rates, while the false negative rates are still relatively high. Interestingly, its performance is fairly constant across the different types of carelessness. As expected, the false negative rate of the traditional methods in general worsens the later the onset of careless responding (while that of CODERS deteriorates the earlier the onset, see Figures 7 and 8 in the main text).
Figure C1: Performance measures for the traditional methods used by Johnson (2005).
@@
Figure C2: False negative rate of the traditional methods used by Johnson (2005) for different types of partially careless respondents.
@@
Decreasing Survey Length
Recall that the changepoint detection method of Shao and Zhang (2010) used in CODERS is based on asymptotic arguments for the number of items being large (see Appendix B). It therefore seems plausible that the ability of CODERS to accurately identify partial carelessness may decline as survey length decreases.
Data Generation
@@
Results
Figure C3 plots the false positive rates for the different survey lengths (in separate columns). As before, CODERS rarely incorrectly flags attentive respondents as careless at any of the considered significance levels, although there is some overrejection for the shortest survey with $p=60$. Figure C4 visualizes the false negative rate (top row) and the mean absolute error of onset estimation (bottom row) for careless respondents across the different surveys lengths (in separate columns). At a given significance level, the false negative rate slightly increases with decreasing survey length compared to the results from the main text. For $p=60$, the false negative rate reaches about 30%–40%, depending on the significance level. In addition, the location of the onset of carelessness is still estimated highly accurately throughout the investigated survey lengths, on average being off by only about one or two items.
Figure C3: False positive rate of CODERS for attentive respondents across decreasing survey length.
@@ Results
Figure C4: Performance measures of CODERS for partially careless respondents across decreasing survey length.
@@
Figure C5: False negative rate of CODERS for different types of partially careless respondents across decreasing survey length.
@@
Figure C6: Mean absolute error of carelessness onset estimation with CODERS for different types of partially careless respondents across decreasing survey length.
@@
Figures C5 and C6 show the false negative rate and the mean absolute error of onset estimation, respectively, grouped by carelessness type (rows) for each considered survey length (columns). In essence, the slight increase in false negative rates seems to be primarily caused by fewer inconsistent careless respondents being detected, whereas the performance for invariable careless responding stays near-perfect. For $p=60$, in particular very few random respondents are correctly identified, but the performance is still acceptable for extreme respondents.
@@
Additional Sample Sizes
@@
Data Generation
We again use the same simulation design of the main text and focus on the baseline onset regime, but we now generate a smaller and a larger number of participants, namely $n=248$ and $n=\text{1,000}$, respectively. Note that the sample size in our simulations should be divisible by 4 since we generate four types of partially careless respondents. This explains the perhaps unusual choice of $n=248$ (rather than $n=250$).
Results
Figures C7C10 visualize the performance measures for $n=248$ and $n=1,000$ simulated participants (in separate columns). The results are similar to those in the main text, with the overall FNR in partially careless respondents increasing only slightly for the smaller sample size and decreasing only slightly for the larger sample size. For random respondents, however, the increase in false negatives in the smaller sample size is notable, but so is the decrease in FNR in the larger sample size (see Figure C9, top row). This can likely be explained by the amount of information the autoencoder has available in order to learn to distinguish between attentive responding and random responding. It is therefore conceivable that the power for detecting random respondents may improve further in even larger sample sizes.
Figure C7: False positive rate of CODERS for attentive respondents across additional sample sizes.
@@
Figure C8: Performance measures of CODERS for partially careless respondents across additional sample sizes.
@@
Figure C9: False negative rate of CODERS for different types of partially careless respondents across additional sample sizes.
@@
Figure C10: Mean absolute error of carelessness onset estimation with CODERS for different types of partially careless respondents across additional sample sizes.
@@
Randomly Choosing from the Middle Answer Categories
@@
Results
Figures C11 and C12 visualizes the performance of CODERS when partially careless respondents are of the middling type. One the one hand, CODERS still almost never incorrectly flags an attentive respondent as partially careless. On the other hand, it identifies very few middling respondents with a false negative rate between about 85% and 95%, depending on the significance level. Hence, these simulations further demonstrate that CODERS requires inconsistent responding to be sufficiently different from attentive responding in order to distinguish between the two.
Figure C11: False positive rate of CODERS for attentive respondents in the presence of partially careless respondents that randomly choose from the middle answer categories.
@@
Figure C12: Performance measures of CODERS for partially careless respondents that randomly choose from the middle answer categories.
@@
In order to further highlight the difficulty of detecting middling carelessness, we again apply the traditional methods personal reliability, psychometric antonym, and the longstring index. The false positive rates and false negative rates are shown in Figure C13. As in the simulation design from the main text, personal reliability flags hardly any respondents, resulting in false positive rates close to 0 and false negative rates close to 1. The performance of the longstring index is similar, which is not surprising because it was designed for detecting straightlining (invariable carelessness) rather than inconsistent responding like middling. Psychometric antonym identifies more middling respondents than CODERS with a false negative rate of about 65%–75%, but it exhibits an inflated false positive rate of up to about 15%.
Figure C13: Performance measures of the traditional methods for partially careless respondents that randomly choose from the middle answer categories.
@@
Results
Figures C14 and C15 visualize the performance of CODERS and Figures C16 and C17 that of the traditional methods. Clearly, CODERS rarely flags any participants, with a false positive rate close to 0 and a false negative rate close to 1 (we therefore omit separate results for the four types of partially careless responding). The same behavior is observed for personal reliability. Psychometric antonym achieves a better false negative rate (about 70%–85%) but struggles with many false positives (up to about 15%). The longstring index yields somewhat better performance, with a relatively low false positive rate and a false negative rate around 65%. As in the simulation design from the main text, it successfully captures straightliners as well as some extreme respondents (see the corresponding section above for a more detailed discussion). In summary, CODERS is simply not designed for detecting more than one changepoint in respondent behavior, such that an extension to multiple changepoint detection may be a fruitful direction for further research.
Figure C14: False positive rate of CODERS for attentive respondents in case of temporary carelessness.
@@
Figure C15: Performance measures of CODERS for partially careless respondents in case of temporary carelessness.
@@
Figure C16: Performance measures for the traditional methods in case of temporary carelessness.
@@
Figure C17: False negative rate of the traditional methods for different types of partially careless respondents in case of temporary carelessness.